OpenStudy (ashontae19):

square root of -200 in standard form

OpenStudy (ashontae19):

@Daddysgirl14

OpenStudy (ashontae19):

@ybarrap

OpenStudy (xapproachesinfinity):

iroot(200)=10iroot(2)

OpenStudy (xapproachesinfinity):

i is imaginary unit

OpenStudy (xapproachesinfinity):

\(\sqrt{-200}=i\sqrt{200}=i10\sqrt{2}\)

OpenStudy (ashontae19):

oh ok i was correct i have no more if u dnt mind

OpenStudy (xapproachesinfinity):

squared by two what do you mean lol

OpenStudy (xapproachesinfinity):

post your other questions in different posts pls!!

OpenStudy (michele_laino):

hint: we can write this: \[\Large - 200 = 200\left\{ {\cos \left( {\pi + 2k\pi } \right) + i\sin \left( {\pi + 2k\pi } \right)} \right\}\]

OpenStudy (xapproachesinfinity):

polar form..

OpenStudy (xapproachesinfinity):

well do the same process you did with the first

OpenStudy (xapproachesinfinity):

-7 you don't have to worry about

OpenStudy (xapproachesinfinity):

just root(-96)

OpenStudy (xapproachesinfinity):

how did root (96) simplify to 4?

OpenStudy (michele_laino):

we have two square roots: \[\Large \begin{gathered} \sqrt { - 200} = \sqrt {200} \left\{ {\cos \left( {\frac{\pi }{2} + k\pi } \right) + i\sin \left( {\frac{\pi }{2} + k\pi } \right)} \right\}, \hfill \\ \hfill \\ k = 0,1 \hfill \\ \end{gathered} \]

OpenStudy (xapproachesinfinity):

root(96)=root(16x6)=4root(6)

OpenStudy (xapproachesinfinity):

oh you have root(6) my bad i miss read your reply

OpenStudy (michele_laino):

\[\Large \begin{gathered} \sqrt { - 96} = \sqrt {96} \left\{ {\cos \left( {\frac{\pi }{2} + k\pi } \right) + i\sin \left( {\frac{\pi }{2} + k\pi } \right)} \right\}, \hfill \\ \hfill \\ k = 0,1 \hfill \\ \end{gathered} \]

OpenStudy (xapproachesinfinity):

your answer is good!

OpenStudy (ashontae19):

yes so i was correct?

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (ashontae19):

thank youu i do not know which one to give medel to since yall both helped me

OpenStudy (xapproachesinfinity):

does not matter :)